OpenStudy (anonymous):

If the length of the heigfht of a trapezoid is 4 meters, the length of one of its bases is 11 meters, and its area is 62 square meters, then what is the measure of the other base?

OpenStudy (anonymous):

area =( h/2) (a+b) 62 = (4/2)(11 + b) 62 = 22 + 2b b = 40/2 = 20 meters